# Personalisation via Dynamic Policy Fusion

### v2

@@ 6 Limitations and Future Work
Even though our proposed dynamic policy fusion approach dynamically optimises two objectives, in its current state, it is restricted to discrete action spaces. Extension to continuous action domains could be explored in future work. Additionally, the choice of hyper-parameters is done using trial and error; reducing or learning the number of parameters will enhance the utility of the method. Our method does not capture the varying preferences of the human. To address this limitation, future research could explore adaptive methods, perhaps based on meta-RL [3] to adjust to changing human preferences with a few shot learning of the LSTM network. The experiments presented in this work also make use of a noise-free ground truth function to simulate human feedback. We acknowledge that in reality, human feedback may be subject to noise, which could affect the robustness of our presented approach. Further research is necessary to adapt our approach to be more robust to noise, the performance of which could then be evaluated with noisy feedback from real humans. Another practical aspect to consider is that the number of trajectories one could expect a human to label is limited. From an initial investigation, our approach does exhibit a drop in performance as fewer labelled trajectories become available. Although this problem could partially be mitigated through the use of semi-supervised methods [29] to automatically generate trajectory score labels, we recognise that more noise-tolerant and feedback-efficient variants of our method need to be explored. Nevertheless, we believe the proposed dynamic policy fusion method presents a novel and an in-principle, practical approach for adapting and personalising already trained policies.
@@ 7 Conclusion
Appendix 0.A Simulating human feedback
@@
Appendix 0.B Training of LSTM
@@
Particulars Value Learning rate $10^{-4}$ Weight decay $10^{-8}$ Gradient clipping $10$ Optimizer ADAM
@@
Particulars Value Learning rate $10^{-3}$ Weight decay $10^{-6}$ Gradient clipping $10$ Optimizer ADAM
@@
Conv layer Input depth Output Depth Kernel size Stride Padding Conv 1 1 16 3 2 0 Conv 2 16 32 3 2 0 Conv 3 32 64 3 1 0 Conv 4 64 128 3 1 0 Conv 5 128 256 3 1 1
@@
ConvT layer Input depth Output Depth Kernel size Stride Padding ConvT 1 256 128 3 1 1 ConvT 2 128 64 3 1 0 ConvT 3 64 32 3 2 0 ConvT 4 32 16 5 1 0 ConvT 5 16 1 4 2 0
@@ Appendix 0.C Training of VAE
Appendix 0.D Different choices of static fusion
@@ 0.D.4 Entropy Weighted
Appendix 0.E Analysis of other choices of policy fusion
@@
where $Z$ is its normalising factor of $\pi_{f}(a|s)$, $Q^{*}=argmax_{a\in\mathcal{A}}Q(s,a)$ and $\zeta=\dfrac{h(Q^{\prime},T_{\psi})}{h(Q,T_{\phi})}$, here $h(Q,T)=\sum_{a}exp\left(\dfrac{Q}{T}\right)$.
Proof
@@
Appendix 0.F MORL baseline Implementation Details
@@
Appendix 0.G Additional experiments
@@ Appendix 0.H DQN Training
Table 11: Table showing the CNN architecture of DQN
Particulars Value Channel 32,64,64 Kernel size 11,5,3 Stride 2,2,1 Padding 0,0,0 FC hidden layer 512
@@
Mode Method Desired region $\uparrow$ Undesired region $\downarrow$ Score $\uparrow$ Preference Static $4.987\pm 0.649$ $0.284\pm 0.045$ Mixed Static $2.372\pm 0.123$ $000.000\pm 0$ 0.000 $0.739\pm 0.014$ Avoidance Static - $000.000\pm 0$ 0.000 $0.974\pm 0.011$
@@
Mode $\alpha$ Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$ Avoidance 0.3 - $16.25\pm 3.45$ $0.15\pm 0.03$ $37.95\pm 0.68$ 0.5 - $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$ 0.7 - $2.75\pm 0.49$ $0.47\pm 0.02$ $33.90\pm 0.91$ Preference 0.3 $9.64\pm 0.79$ $0.15\pm 0.02$ $37.58\pm 0.60$ 0.5 $11.58\pm 1.12$ $0.07\pm 0.06$ $38.14\pm 0.30$ 0.7 $23.74\pm 1.58$ $0.09\pm 0.01$ $38.42\pm 0.49$ Mixed 0.3 $11.90\pm 0.93$ $11.45\pm 1.27$ $0.13\pm 0.02$ $38.03\pm 0.55$ 0.5 $12.61\pm 1.20$ $12.36\pm 1.58$ $0.07\pm 0.01$ $36.97\pm 0.59$ 0.7 $23.18\pm 2.55$ $6.13\pm 1.47$ $0.11\pm 0.03$ $37.64\pm 0.51$

### v3

@@ 6 Limitations and Future Work
Even though our proposed dynamic policy fusion approach dynamically optimises two objectives, in its current state, it is restricted to discrete action spaces. Extension to continuous action domains could be explored in future work. Additionally, the choice of hyper-parameters is done using trial and error; reducing or learning the number of parameters will enhance the utility of the method. Our method does not capture the varying preferences of the human. To address this limitation, future research could explore adaptive methods, perhaps based on meta-RL [3] to adjust to changing human preferences with a few shot learning of the LSTM network. The experiments presented in this work also make use of a noise-free ground truth function to simulate human feedback. We acknowledge that human feedback may be subject to noise, which could affect the robustness of our presented approach. Further research is necessary to adapt our approach to be more robust to noise, the performance of which could then be evaluated with noisy feedback from real humans. Another practical aspect to consider is that the number of trajectories one could expect a human to label is limited. From an initial investigation, our approach does exhibit a drop in performance as fewer labelled trajectories become available. Although this problem could partially be mitigated through the use of semi-supervised methods [30] to automatically generate trajectory score labels, we recognise that more noise-tolerant and feedback-efficient variants of our method need to be explored. Nevertheless, we believe the proposed dynamic policy fusion method presents a novel and an in-principle, practical approach for adapting and personalising already trained policies.
@@ 7 Conclusion
8 Statements and Declarations
Funding The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.
Competing Interests The authors have no relevant financial or non-financial interests to disclose.
Author Contributions All authors contributed to the study conception, brainstorming and formulating the idea. Material preparation, data collection, experimentation and analysis were performed by Ajsal Shereef Palattuparambil. The first draft of the manuscript was written by Ajsal Shereef Palattuparambil, and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
Ethics Approval and Consent to Participate Research does not involve Human Participants and/or Animals.
Consent for Publication Yes.
data availability statement The code used for the experimentation and analysis is available here.
Appendix A Simulating human feedback
@@
Appendix B Training of LSTM
@@
Particulars Value
Learning rate $10^{-4}$
Weight decay $10^{-8}$
Gradient clipping $10$
Optimizer ADAM
@@
Particulars Value
Learning rate $10^{-3}$
Weight decay $10^{-6}$
Gradient clipping $10$
Optimizer ADAM
@@
Conv layer Input depth Output Depth Kernel size Stride Padding
Conv 1 1 16 3 2 0
Conv 2 16 32 3 2 0
Conv 3 32 64 3 1 0
Conv 4 64 128 3 1 0
Conv 5 128 256 3 1 1
@@
ConvT layer Input depth Output Depth Kernel size Stride Padding
ConvT 1 256 128 3 1 1
ConvT 2 128 64 3 1 0
ConvT 3 64 32 3 2 0
ConvT 4 32 16 5 1 0
ConvT 5 16 1 4 2 0
@@ Appendix C Training of VAE
Appendix D Different choices of static fusion
@@ D.4 Entropy Weighted
Appendix E Analysis of other choices of policy fusion
@@
where $Z$ is its normalising factor of $\pi_{f}(a|s)$, $Q^{*}=max_{a\in\mathcal{A}}Q(s,a)$ and $\zeta=\dfrac{h(Q^{\prime},T_{\psi})}{h(Q,T_{\phi})}$, here $h(Q,T)=\sum_{a}exp\left(\dfrac{Q}{T}\right)$.
Proof.
@@
Appendix F MORL baseline Implementation Details
@@
Appendix G Additional experiments
@@ Appendix H DQN Training
Table 11: Table showing the CNN architecture of DQN
Particulars Value
Channel 32,64,64
Kernel size 11,5,3
Stride 2,2,1
Padding 0,0,0
FC hidden layer 512
@@
Mode Method Desired region $\uparrow$ Undesired region $\downarrow$ Score $\uparrow$
Preference Static $4.987\pm 0.649$ $0.284\pm 0.045$
Mixed Static $2.372\pm 0.123$ $000.000\pm 0$ 0.000 $0.739\pm 0.014$
Avoidance Static - $000.000\pm 0$ 0.000 $0.974\pm 0.011$
@@
Mode $\alpha$ Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$
Avoidance 0.3 - $16.25\pm 3.45$ $0.15\pm 0.03$ $37.95\pm 0.68$
0.5 - $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$
0.7 - $2.75\pm 0.49$ $0.47\pm 0.02$ $33.90\pm 0.91$
Preference 0.3 $9.64\pm 0.79$ $0.15\pm 0.02$ $37.58\pm 0.60$
0.5 $11.58\pm 1.12$ $0.07\pm 0.06$ $38.14\pm 0.30$
0.7 $23.74\pm 1.58$ $0.09\pm 0.01$ $38.42\pm 0.49$
Mixed 0.3 $11.90\pm 0.93$ $11.45\pm 1.27$ $0.13\pm 0.02$ $38.03\pm 0.55$
0.5 $12.61\pm 1.20$ $12.36\pm 1.58$ $0.07\pm 0.01$ $36.97\pm 0.59$
0.7 $23.18\pm 2.55$ $6.13\pm 1.47$ $0.11\pm 0.03$ $37.64\pm 0.51$